Long Term Shifts in Radio Content

Matthew Gray

February 27, 1997

Introduction

Radio, over the years has shifted from being an interactive and experimental medium, to being a broadcast medium. During the time which radio has been a primarily broadcast, it has gone through two substantial phases. These two stages are the broadcast narrative and broadcast music stages. This paper will focus on this latter shift.

Radio’s initial broadcast content focused on narrative content. Radio shows would tell episodic stories, not too distant in character from present day television broadcasts. In fact, it seems clear that the popularization of television is what brought about this shift in content.

Narrative Content

While in the earlier phase of radio, there was a mix of both musical and narrative programming, the narrative content was typically awkward in format, though a variety of clever measures to overcome this awkwardness were adopted.

It is not surprising that this format was somewhat difficult to adapt to. Essentially all narrative forms before had some visual component, or, in the case of pure text, were random access. Even simple traditional oral storytelling came with the features of gesture, movement, and facial nuance. Clearly radio lacked these, and more obviously lacked the even more directly visual features of theatre and similar narrative forms.

Radio narratives worked around these problems by incorporating extensive use of verbal description of visual information by the characters (“Look at that, isn’t that a pretty shiny stone in that gold ring”). While strained, it served the purpose. Many narratives also succeeded by focusing disproportionately on dialog, or other components that can be effectively expressed almost completely aurally.

While radio did succeed in telling stories, it lacked much. One might try to come to radio’s defense as a valuable narrative form, but given that upon the arrival of television, radio narrative essentially disappeared entirely from radio, the case would be somewhat weak. Radio managed as a functional, but inferior narrative form.

Non-narrative Content

With the rise of television, radio became entire displaced as a narrative medium and radio was forced to find other types of content to focus on. Capitalizing on its fundamentally aural nature, music programming became the dominant content of radio.

Radio today continues to be dominated by music programming, but finds a range of other content that plays to the features and restrictions of radio. “Talk radio” is well suited to this medium, as the complete focus is on conversation. While this format works on televesion as well, it adds little, and hence radio can compete in this area. Other content, often music related, is provided by DJs as an add on to the music content.

Further, the advent of the car radio, and other portable radios opened up the demand for news, traffic and weather radio. In this case the incompleteness of radio becomes an advantage. A more complete medium, like television demands sufficiently more attention that it is not suitable for use in a car, or other on-the-go activities.

Speculation

It is worth considering what the impact of Internet technologies is likely to have on radio. It doesn’t seem that the Internet is likely to usurp any substantial amount of radio’s current content, the way television did of radio’s narrative content, but there are certainly possibilities. Radio style audio broadcasting is now possible over the Internet, as well as distribution of audio data via other means. While having many of the same features of radio, Internet audio capabilities remain sufficiently different in terms of capabilities and limitations that it is not likely to be an area of transition.

Equally so, news and talk radio content is less likely to be adopted by the Internet in the same way that radio presents them. Clearly, news and weather, and to some extent traffic information is available on the Internet, but due to the way it is accessed, it is not likely to compete with radio any time soon. Equally so, discussion and “talk” systems on the Internet are more likely to become their own media format altogether, rather than displace talk radio.

Conclusion

Radio has been around long enough to be considered a mature medium. One of the important steps in this maturation was the transition from narrative content, which did not suit the medium well, to radio specific forms of non-narrative content.

The introduction of “new media” doesn’t seem likely to substantially shift radio’s current content. How this new media might affect the content of television may in fact be more similar to the way television affected radio.